OpenStudy (anonymous):

8- 40x =13 keep in fraction form

OpenStudy (lilg132):

-40x = 13 -8 -40x = 5 x = -0.125

OpenStudy (anonymous):

okay but keep it in a fraction and why does it become negative?

OpenStudy (anonymous):

8-13=40x -5=40x \[x=\frac{-5}{40}=\frac{-1}{8}\]

OpenStudy (lilg132):

-1/8

OpenStudy (lilg132):

sumbul where can i find the fractions sign in equations?

OpenStudy (anonymous):

write it in this format click on equation button then write frac{1}{8} etc you will get \[\frac{1}{8}\] write numerator in first bracket and denominator in second bracket

OpenStudy (anonymous):

fraction sign is not available in equation but you can use above method to write in fractions

OpenStudy (anonymous):

x=-\frac{1}{8}\[x=-\frac{1}{8} \]

OpenStudy (anonymous):

hey